OpenStudy (cindy):

simplify 2(1-d) - (2d+1) (2-2d) (2d+1) 2-4d+1 = 3-4d is that correct

OpenStudy (immanuelv):

no....

OpenStudy (immanuelv):

its 1-4d

OpenStudy (cindy):

how...can you show me thesteps

OpenStudy (immanuelv):

2(1-d) - (2d+1) (2-2d) + (-2d-1) 2-4d-1 1-4d

OpenStudy (immanuelv):

is that helpful? i think u forgot to change the sign!

OpenStudy (cindy):

yeah thats helpfull..thanks

OpenStudy (immanuelv):

do u have any other questions?

OpenStudy (cindy):

yeah i do...can yo help me on this one its kinda hed simplify 1+ 1/2+1/3

OpenStudy (immanuelv):

so first you need common denominator....(bottom value) which will be 6 (2*3) so : (6/6)+ (3/6)+ (2/6) now... add the numerators : (Top values) it will give you (11/6) the answer is (11/6)

OpenStudy (cindy):

im sorry but i stil dnt understand

OpenStudy (immanuelv):

you want all the bottom values to be the same.... therefore you multiply all of them together to get a common denominator... and it is 6 so when you add them.... it will be (6/6) = 1, (3/6)= (1/2) , and (2/6)= (1/3) so you add 3+6+2 = 11 then put it over the common denominator =6

OpenStudy (cindy):

oka but the equation is 1+ a fration witch is 1/ andthe wholebotom equatio is 2+1/3

OpenStudy (immanuelv):

then it is 10/7

OpenStudy (immanuelv):

is the equation (1)+ (1/(2+(1/3)))

OpenStudy (cindy):

take alook

OpenStudy (immanuelv):

that's what i wrote. 1+ (1)/(2)+(1/3) now 2+ (1/3) = 7/3 when you divide 1 by 7/3 it flips: therefore it becomes 3/7 now the equation is 1+ 3/7 common denominator= 7 7/7 + 3/7 = 10/7

OpenStudy (immanuelv):

do u understand?

OpenStudy (cindy):

yes now i do thankyou much

OpenStudy (immanuelv):

no prob..... glad to help you!